Density and Percent Composition Graph

Sugar Solution Percents (x-axis): 0.00%, 4.00%, 8.00%, 12.00%, 16.00%

Density of Solutions (y-axis): 0.998141 (0%), 1.0080 (4%), 1.0232 (8%), 1.0412 (12%), 1.0532 (16%)

I just want to know how you would go about properly organizing the y-axis with all the densities
Answer #1

Hi,

To prepare the graph: sugar solution (%) vs. solution density (%), you must choose a scale in which all the points enter.

For the scale of the "Y" axis, 7 values were taken that vary from 0.01 to 0.01. While on the axis of the "X" 5 values were taken in which they vary from 5 to 5.
Subsequently, each value was intersected until the points that drew a slope were obtained.

If you find it difficult to locate decimal numbers, I suggest you do the following,
To choose the scale you will take the highest value of your data and subtract it less the least value, then divide it by the amount of values:



A value of 0.01 was obtained, then the scale of "Y" will be made of 0.01 in 0.01.

You will add to your lowest value of your data (0.99) 0.01 in 0.01, every 0.01 added will be the value of your Y axis, until you reach your last value.
1st value: 0.99
2nd value: 0.99 + 0.01 = 1
3rd value: 1+ 0.01 = 1.01
4th value: 1.01 + 0.01 = 1.02
5th value: 1.02 + 0.01 = 1.03
6th value: 1.03 + 0.01 = 1.04
7th value: 1.04 + 0.01 = 1.05
8th value: 1.05 + 0.01 = 1.06




If you liked the resolution, a thumbs up would help me a lot. Thank you very much and success. Regards